# Quantization Meets Reasoning: Exploring and Mitigating Degradation of Low-Bit LLMs in Mathematical Reasoning

### v3

@@ 1 Introduction
Guided by this view, as shown on Figure 1, we adopt a lightweight measure–locate–restore loop that operates directly on the quantized model. We first locate the initial erroneous step, then apply small-scale supervised/preference tuning on a compact “Silver Bullet” set designed to target the observed weaknesses. In our experiments, fine-tuning on as few as 332 curated examples for 3–5 minutes on a single GPU is sufficient to recover the mathematical reasoning accuracy of W4A16 models toward their full-precision baselines, while preserving PTQ’s memory and latency benefits.
@@ 2.2 Reasoning Ability Optimization in Large Language Models
Orthogonal to prompting, alignment research pursues systematic post-training refinements. Instruction tuning on diversified task mixtures (FLAN) (Wei et al., 2021) and lightweight data-curation pipelines (Alpaca) (Taori et al., 2023) make models broadly helpful, while Direct Preference Optimization (DPO) (Rafailov et al., 2024) offers sample-efficient preference learning without full RLHF. Reliability has been pushed along two complementary axes: self-consistency voting for answer selection (Wang et al., 2022) and process-level supervision with stepwise reward models (Lightman et al., 2023b), both grounded in verifiable-reasoning theory (Creswell & Shanahan, 2022).
@@ 3.1.3 Detailed Examination of Reasoning Process
Qualitative Analysis To systematically investigate the underlying reasons for degradation in quantized models, we performed a qualitative error analysis inspired by established categorizations from previous literature (Brown et al., 2016), (Delastri & Lolang, 2023) and (Kurudirek et al., 2023), which categorize real world student errors in mathematical problem solving. Building on these frameworks, we conduct a qualitative analysis by classifying model-generated errors into seven fine-grained error types, organized under four high-level categories. The definitions of these error types are detailed as follows:
@@
Table 1: Comparison of quantization methods applied to the Llama-3 and Qwen2.5 model families. AWQ and GPTQ employ 4-bit weight and 16-bit activation quantization, whereas SmoothQuant uses 8-bit weight and 8-bit activation quantization.
@@
Quantitative Analysis and Error Assessment Pipeline
@@
1. Expert Model Judgement: For each instance in which a quantized model produces an incorrect answer, we employ a dedicated “expert model” to analyze the error. The expert model is tasked with: (a) identifying the first occurrence of an error, (b) specifying the exact step where the error is introduced, (c) assigning an error category based on a nested classification scheme, and (d) providing an explanation along with a confidence score for its determination.
2. Majority Voting: To curb hallucinations and improve evaluation reliability, we apply a three-stage majority-vote protocol to the outputs of five language models—DeepSeek-R1 (Guo et al., 2025) (primary), GPT-4o, GPT-4, Qwen-Max (Yang et al., 2025), and DeepSeek-V3 (Liu et al., 2024). Instances of disagreement are flagged for further review, ensuring consistency and minimizing spurious judgments. Rule1-Unanimous agreement: If all four auxiliary models concur with the reference judgment from DeepSeek-R1, the answer is accepted. Rule2-Simple majority: If exactly three auxiliary models concur with DeepSeek-R1, the answer is likewise accepted. Rule3-Escalation: Otherwise, the instance is forwarded to two independent human annotators for arbitration.
3. Human Annotation: For cases with conflicting assessments from the majority vote, we introduce two human annotators to manually review is conducted. The annotator need to follow the annotation document and review the explanations of five expert models then give the final assessment. Additionally, we also randomly sample 2% of the passed evaluated cases to verify the accuracy and consistency of the automated judgments. The annotation documents are detailed in Appendix C.
This pipeline generally aligns with human error analysis and minimizes misclassifications and inconsistencies. However, we also encountered unexpected interesting scenarios. For instance, when the canonical answer is $\frac${11}{2}" but the quantized model outputs "5.5" the expert model sometimes erroneously concludes that there is "No Error" due to subtle discrepancies in reasoning or formatting even if we do not give this type of judgment. Such findings underscore both the robustness and trustiness of our judge framework.
@@ 4.1 Evaluating Quantization Effects
Smaller-scale models suffer more severe losses in complex reasoning ability after quantization: Across all quantization methods, smaller-scale models consistently demonstrate increased vulnerability to quantization. Specifically, the Qwen2.5-0.5B-Instruct model experiences accuracy drops exceeding 60% post-quantization, whereas the larger Qwen2.5-7B-Instruct model incurs only a modest degradation of approximately 2–3%. This trend is also corroborated within the Llama3 model series. To rule out potential biases arising from larger models more readily fitting the calibration datasets, we further validated our findings using calibration datasets of varying sizes, consistently obtaining similar results. This evidence suggests that smaller models are more adversely affected by quantization-induced shifts in feature distributions, thereby experiencing more severe performance declines in complex mathematical reasoning tasks.
Performance degradation becomes more pronounced as the task complexity increases: We evaluated model accuracy across three mathematical reasoning benchmarks of varying difficulty levels. Our results indicate a clear trend wherein performance degradation exacerbates as task complexity rises. Among these, AIME represents the most challenging benchmark, with even full-precision models constrained by their scale unable to solve all problems effectively. The MATH dataset, characterized by evenly distributed difficulty tiers, poses intermediate-level complexity, while GSM8K is comparatively less challenging. Notably, quantized models exhibited relatively minor accuracy losses on the simpler GSM8K benchmark, with an average performance decline of only 7.16%. In contrast, the MATH dataset incurred a more pronounced average degradation of 15.18%. The most severe impact was observed on the highly challenging AIME benchmark, where quantization frequently led to complete failure in problem-solving capability.
@@ 4.2 Error Taxonomy and Its Shift Under Quantization
A careful case study shows that the drop in Reasoning Errors is largely an artefact of our evaluation protocol. Reasoning-type failures typically emerge late in a solution, but the pipeline records only the first erroneous step. Quantization pushes simpler mistakes to earlier stages, so these early slips mask any subsequent logical flaws. In many instances the overall reasoning chain remains valid, yet a premature numerical inaccuracy still leads to an incorrect final answer. Absolute accuracies and additional illustrative examples are reported in Appendix D.
@@ 4.5 Discussion
Fine-tuning on the compact failure-targeted split substantially restores performance on MATH500 while also boosting GSM8K, and does so without hurting broad-domain reasoning as measured by MMLU. This confirms that a small, carefully constructed dataset can achieve capability recovery rather than memorization, aligning with our design goal of lightweight in-situ correction.
@@
Ablation results (Table 2) demonstrate that regenerating solutions from the first erroneous step onward yields larger gains than either (i) adopting the full-precision model’s complete derivation (All-NS), or (ii) training on size-matched random problems (Rand-NS). Focusing supervision precisely at the point of failure proves to be the most effective strategy for restoring mathematical reasoning.
@@ 5 Conclusion
In this study, we present a systematic study of quantization-induced degradation in the mathematical reasoning abilities of large language models, revealing that low-bit post-training quantization especially harms smaller models’ procedural and execution skills. To address this, we propose a lightweight recovery pipeline that combines step-aligned error analysis with targeted fine-tuning on compact, automatically constructed “Silver Bullet” datasets. Experiments show that, with minimal data and compute, quantized models can recover reasoning performance to match their full-precision counterparts while preserving efficiency and general capabilities. Our approach offers a practical and extensible solution for deploying quantized LLMs in resource-constrained settings, and opens avenues for robust reasoning restoration in broader domains.
@@ Appendix E Experiment Results
E.1 Case Statistics
Detailed statistics of all error types. The total number of cases varies slightly across models due to differences in error rates and scores.

### v4

@@ 1 Introduction
Guided by this view, as shown on Figure 1, we adopt a lightweight measure–locate–restore loop that operates directly on the quantized model. We first locate the initial erroneous step, then apply small-scale supervised/preference tuning on a compact "Silver Bullet" set designed to target the observed weaknesses. In our experiments, fine-tuning on as few as 332 curated examples for 3–5 minutes on a single GPU is sufficient to recover the mathematical reasoning accuracy of W4A16 models toward their full-precision baselines, while preserving PTQ’s memory and latency benefits.
@@ 2.2 Reasoning Ability Optimization in Large Language Models
Orthogonal to prompting, alignment research pursues systematic post-training refinements. Instruction tuning on diversified task mixtures (FLAN) (Wei et al., 2021) and lightweight data-curation pipelines (Alpaca) (Taori et al., 2023) make models broadly helpful, while Direct Preference Optimization (DPO) (Rafailov et al., 2024) offers sample-efficient preference learning without full RLHF. Reliability has been pushed along two complementary axes: self-consistency voting for answer selection (Wang et al., 2022) and process-level supervision with stepwise reward models (Lightman et al., 2023b), both grounded in verifiable-reasoning theory (Creswell and Shanahan, 2022).
@@ 3.1.3 Detailed Examination of Reasoning Process
Qualitative Analysis.
To systematically investigate the underlying reasons for degradation in quantized models, we performed a qualitative error analysis inspired by established categorizations from previous literature (Brown et al., 2016), (Delastri and Lolang, 2023) and (Kurudirek et al., 2023), which categorize real world student errors in mathematical problem solving. Building on these frameworks, we conduct a qualitative analysis by classifying model-generated errors into seven fine-grained error types, organized under four high-level categories. The definitions of these error types are detailed as follows:
@@
Table 1: Comparison of quantization methods applied to the Llama-3 and Qwen2.5 model families. AWQ and GPTQ employ 4-bit weight and 16-bit activation quantization, whereas SQ (SmoothQuant) uses 8-bit weight and 8-bit activation quantization.
@@
Notes: Van. denotes the vanilla full-precision baseline; Inst. is an abbreviation of Instruct.
Quantitative Analysis and Error Assessment Pipeline.
@@
1. Expert Model Judgement
: For each instance in which a quantized model produces an incorrect answer, we employ a dedicated "expert model" to analyze the error. The expert model is tasked with: (a) identifying the first occurrence of an error, (b) specifying the exact step where the error is introduced, (c) assigning an error category based on a nested classification scheme, and (d) providing an explanation along with a confidence score for its determination.
2. Majority Voting
: To curb hallucinations and improve evaluation reliability, we apply a three-stage majority-vote protocol to the outputs of five language models—DeepSeek-R1 (Guo et al., 2025) (primary), GPT-4o, GPT-4, Qwen-Max (Yang et al., 2025), and DeepSeek-V3 (Liu et al., 2024). Instances of disagreement are flagged for further review, ensuring consistency and minimizing spurious judgments. Rule1-Unanimous agreement: If all four auxiliary models concur with the reference judgment from DeepSeek-R1, the answer is accepted. Rule2-Simple majority: If exactly three auxiliary models concur with DeepSeek-R1, the answer is likewise accepted. Rule3-Escalation: Otherwise, the instance is forwarded to two independent human annotators for arbitration.
3. Human Annotation
: For cases with conflicting assessments from the majority vote, we introduce two human annotators to manually review is conducted. The annotator need to follow the annotation document and review the explanations of five expert models then give the final assessment. Additionally, we also randomly sample 2% of the passed evaluated cases to verify the accuracy and consistency of the automated judgments. The annotation documents are detailed in Appendix C.
This pipeline is intentionally designed to be conservative and to avoid spurious "false consensus" in the automatic labels. Among several strong candidate judges, we select DeepSeek-R1 as the pivot because its reflective, stepwise chain-of-thought makes it particularly suitable for localizing the first erroneous step and producing structured explanations; in a small pilot on a random subset of failures, its error-type predictions also showed the highest agreement with two human annotators. For each incorrect model output, we then collect judgments from all five expert models and accept an automatic label only when at least three of the four auxiliary models concur with the pivot, that is, at least four out of five models agree on both the error type and its location; otherwise the instance is escalated to human annotators. In addition, we randomly sample two percent of the automatically accepted cases for manual audit.
Under this protocol, our automated error-assessment pipeline matches the final human judgment on 97.2% of 9,908 failure cases, with the remaining discrepancies concentrated in borderline situations, for example when the canonical answer is $\frac${11}{2}" but the quantized model outputs the numerically equivalent "5.5" and some expert models still flag an error due to subtle reasoning or formatting differences. These observations suggest that our judge framework achieves high precision at the cost of slightly lower recall, which is appropriate for the downstream analyses in this paper; further implementation details and annotation guidelines are provided in Appendix C.
@@ 4.1 Evaluating Quantization Effects
Smaller-scale models suffer more severe losses in complex reasoning ability after quantization:
Across all quantization methods, smaller-scale models consistently demonstrate increased vulnerability to quantization. Specifically, the Qwen2.5-0.5B-Instruct model experiences accuracy drops exceeding 60% post-quantization, whereas the larger Qwen2.5-7B-Instruct model incurs only a modest degradation of approximately 2–3%. This trend is also corroborated within the Llama3 model series. To rule out potential biases arising from larger models more readily fitting the calibration datasets, we further validated our findings using calibration datasets of varying sizes, consistently obtaining similar results. This evidence suggests that smaller models are more adversely affected by quantization-induced shifts in feature distributions, thereby experiencing more severe performance declines in complex mathematical reasoning tasks.
Performance degradation becomes more pronounced as the task complexity increases:
We evaluated model accuracy across three mathematical reasoning benchmarks of varying difficulty levels. Our results indicate a clear trend wherein performance degradation exacerbates as task complexity rises. Among these, AIME represents the most challenging benchmark, with even full-precision models constrained by their scale unable to solve all problems effectively. The MATH dataset, characterized by evenly distributed difficulty tiers, poses intermediate-level complexity, while GSM8K is comparatively less challenging. Notably, quantized models exhibited relatively minor accuracy losses on the simpler GSM8K benchmark, with an average performance decline of only 7.16%. In contrast, the MATH dataset incurred a more pronounced average degradation of 15.18%. The most severe impact was observed on the highly challenging AIME benchmark, where quantization frequently led to complete failure in problem-solving capability.
We also evaluate the quantization-induced degradation on both thinking-mode models and larger-scale models for mathematical reasoning tasks, the detailed results are reported in Appendix E.1. Together with the corresponding evaluations under the same quantization settings on general-purpose benchmarks, the detailed results are reported in Appendix E.2.
@@ 4.2 Error Taxonomy and Its Shift Under Quantization
The apparent reduction in Reasoning Errors after quantization arises from a statistical masking effect induced by our standardized evaluation protocol. Each trajectory is assigned a single error type according to the first erroneous step. In contrast, reasoning type failures in our taxonomy usually occur later in the solution when global logical consistency or boundary conditions are evaluated. Quantization increases the likelihood of earlier and simpler mistakes such as Conceptual or Execution Errors, and these early failures "hide" subsequent reasoning flaws from the statistics. Our case study confirms that many trajectories labeled as other error types still contain additional reasoning problems at later steps, although these later issues are not recorded because they occur after the first mistake. This masking effect therefore reflects how quantization reshapes the distribution of observed first errors under a reproducible and unambiguous protocol, not an actual disappearance of deeper reasoning mistakes. Absolute accuracies and examples are provided in Appendix D.
@@ 4.5 Discussion
Fine-tuning on the compact failure-targeted split substantially restores performance on MATH500 while also boosting GSM8K, and does so without hurting broad-domain reasoning as measured by MMLU. This intervention uses only a few hundred preference pairs and a few minutes of training on a single GPU, yet closes much of the gap between quantized and full-precision models. This confirms that the "Silver Bullet" provides sample-efficient capability recovery rather than simple memorization.
@@
Ablation results in Table 2 is run under a strictly matched data and compute budget: each setting uses the same number of stepwise preference pairs and identical training hyperparameters, and only the selection of problems and traces is varied. Under this controlled setup, training on our error-targeted stepwise All split consistently outperforms both size-matched random supervision (Random) and the non-stepwise variant (Non-Step) that adopts full-precision derivations without restarting from the first erroneous step. On average, All improves accuracy by about 1.7 points over Random and by about 5.6 points over Non-Step, with the largest gap of 6.2 points on MATH500 for the most quantization-sensitive configuration (Qwen2.5-3B with GPTQ). These results indicate that locating the earliest quantization-induced failure and regenerating the remaining steps from that point provides a much stronger learning signal than unconditioned math supervision or full-solution positives, especially in the low-data regime we consider.
@@ 5 Conclusion
In this study, we present a systematic study of quantization-induced degradation in the mathematical reasoning abilities of large language models, revealing that low-bit post-training quantization especially harms smaller models’ procedural and execution skills. To address this, we propose a lightweight recovery pipeline that combines step-aligned error analysis with targeted fine-tuning on compact, automatically constructed "Silver Bullet" datasets. Experiments show that, with minimal data and compute, quantized models can recover reasoning performance to match their full-precision counterparts while preserving efficiency and general capabilities. Our approach offers a practical and extensible solution for deploying quantized LLMs in resource-constrained settings, and opens avenues for robust reasoning restoration in broader domains.
@@ Appendix E Experiment Results
E.1 Quantization on Larger Models
To strengthen the generality of our conclusions and bringing potential insights, we have conducted additional experiments on the Qwen3 series, which natively support a "thinking mode" as the default configuration. These models already integrate internal CoT-like mechanisms, making them representative of both standard and "thinking" (CoT) LLMs. We evaluated Qwen3-8B, Qwen3-14B, and Qwen3-32B under identical quantization settings (bit-width, calibration set, and hyperparameters) as other qutization methods.
Table 3: Performance of Qwen3 models on MATH under different quantization methods.
Qwen3-8B Qwen3-14B Qwen3-32B
Van. AWQ GPTQ Van. AWQ GPTQ Van. AWQ GPTQ
MATH 55.88 53.96 53.58 63.52 62.28 63.12 66.92 65.72 65.94
Note: Van. denotes the vanilla full-precision baseline.
These results yield several key insights:
  • Larger models show greater quantization robustness. Accuracy degradation from full-precision to quantized versions diminishes significantly as model size increases, consistent with observations from other studies. We believe because larger models possess a richer parameter space, which grants stronger robustness to quantization-induced numerical errors when mapping from high-precision to low-precision data formats.
  • Mild regularization effects appear in quantized models. You can find an interesting result that Qwen3-14B under GPTQ slightly outperforms its vanilla counterpart, suggesting that moderate compression may enhance generalization, a phenomenon also noted in recent quantization studies. Notably, this effect has also been frequently cited as one of the underlying reasons why Quantization-Aware Training (QAT) can sometimes improve the generalization ability of post-quantized models.
  • Scalability of the "Silver Bullet" principle. The consistently small degradation across larger models supports our hypothesis that targeted recovery using compact, well-curated data is even more effective for high-capacity models with stronger learning abilities.
E.2 Additional Evaluation on General-Reasoning Benchmarks
To complement our analysis on mathematical reasoning, we additionally evaluate Qwen2.5-0.5B/1.5B/3B/7B-Instruct models on several widely used benchmarks that probe general science question answering, commonsense reasoning, and instruction following:
  • ARC Easy and ARC Challenge: a science exam multiple-choice benchmark (grades 3–9) with two difficulty splits. Most questions provide four options and the Challenge split requires more complex reasoning (Clark et al., 2018).
  • HellaSwag: a commonsense inference benchmark with 70k multiple-choice questions. Each item provides a scenario and four possible continuations; the distractors are adversarially generated to fool models while remaining trivial for humans (Zellers et al., 2019).
  • IFEval: an instruction-following benchmark built from verifiable constraints (for example “write more than 400 words”) that focuses on controllable instructions and reduces the bias of LLM-based judges (Zhou et al., 2023).
  • CommonSenseQA: a 12k-question multiple-choice benchmark that requires various forms of commonsense knowledge, with one correct answer and four distractors per question (Talmor et al., 2019).
For each benchmark we compare the vanilla full-precision model with its AWQ and GPTQ quantized counterparts, and summarize the average degradation in Table 4. Compared with the larger drops on mathematical reasoning tasks, the performance drops on general commonsense and language-understanding benchmarks are substantially smaller, both in absolute score reduction and in relative percentage. This supports our claim that post-training quantization disproportionately affects mathematical reasoning ability while having only mild impact on general language capabilities.
Table 4: Average accuracy degradation of Qwen2.5-Instruct models on additional benchmarks under post-training quantization. Negative values indicate performance drop compared with the corresponding full-precision models.
Benchmark Type Avg. accuracy drop $\downarrow$ (points) Avg. relative drop $\downarrow$ (%)
ARC-c General science QA -3.56 -4.10
ARC-e General science QA -3.59 -4.76
CommonSenseQA Commonsense QA -3.38 -5.06
HellaSwag Commonsense -2.35 -3.95
IFEval Instruction following -2.61 -5.01
GSM8K Grade-school math -6.78 -13.21
MATH Competition math -10.56 -29.84
The detailed per-model results are listed in Table 5. We report accuracy for each Qwen2.5-Instruct checkpoint and for each quantization method.
Table 5: Accuracy (%) of Qwen2.5-Instruct models on general benchmarks before and after quantization.
Model Scale Method ARC-c ARC-e CommonSenseQA HellaSwag IFEval
0.5B Vanilla 46.44 65.43 59.38 39.43 36.69
AWQ 51.86 64.90 53.81 37.81 36.81
GPTQ 43.39 50.97 50.37 36.47 32.01
1.5B Vanilla 77.97 89.95 76.00 62.19 50.96
AWQ 72.54 84.13 73.38 58.62 47.60
GPTQ 71.86 86.42 71.42 60.19 46.04
3B Vanilla 84.75 91.53 78.62 76.62 68.23
AWQ 75.59 88.36 76.41 73.28 65.47
GPTQ 78.64 89.77 77.56 73.52 64.87
7B Vanilla 86.10 92.59 84.19 85.18 77.82
AWQ 86.44 93.47 82.47 84.38 76.86
GPTQ 84.75 92.24 83.95 83.75 76.86
Note: Inst. is an abbreviation of Instruct.
@@ E.2 Additional Evaluation on General-Reasoning Benchmarks
Notes: Van. denotes the vanilla full-precision baseline; Inst. is an abbreviation of Instruct.
E.3 Case Statistics
Table 6 shows detailed statistics of all error types. The total number of cases varies slightly across models due to differences in error rates and scores.
E.4 Subject-wise and Difficulty-wise Degradation on MATH
To better understand how quantization-induced degradation relates to problem structure, we leverage the rich annotations in the MATH dataset, which covers multiple subject domains (such as algebra, geometry, number theory and combinatorics) and five difficulty levels (Level 1 to Level 5). For each model scale and quantization method, we compute the distribution of errors across mathematical subfields and difficulty levels. This analysis connects quantization sensitivity with different types of reasoning and provides practical guidance for deploying quantized models under varying reasoning complexities.
Table 7: Distribution of errors across mathematical domains on MATH under different quantization settings. All values are percentages. All experiments are conducted on Qwen2.5-Instruct models.
Model Method Number Theory Counting & Prob. Interm. Algebra Algebra Geometry Prealgebra Precalculus
0.5B Van. 11.59 10.33 20.48 18.65 10.30 16.02 12.63
AWQ 11.26 9.82 18.77 22.24 9.84 16.68 11.39
GPTQ 11.08 9.80 18.74 22.57 9.65 16.71 11.45
1.5B Van. 11.31 10.67 22.98 15.79 10.67 13.73 14.86
AWQ 11.39 10.26 20.16 19.16 10.62 15.72 12.68
GPTQ 11.57 10.28 21.74 18.99 10.28 13.94 13.56
3B Van. 10.40 10.35 24.85 12.91 11.94 12.07 17.49
AWQ 10.74 10.78 25.05 13.92 11.13 12.29 16.09
GPTQ 11.23 9.98 24.49 14.58 11.20 12.30 16.22
7B Van. 7.96 10.31 26.28 11.75 12.87 11.91 18.91
AWQ 9.58 9.88 25.89 11.87 12.22 11.67 18.90
GPTQ 8.56 10.50 27.13 11.20 12.39 11.60 18.62
Note: Van. denotes the vanilla full-precision baseline.
Table 8: Distribution of errors across difficulty levels on MATH under different quantization settings. All values are percentages.
Model Level 1 Level 2 Level 3 Level 4 Level 5
Qwen2.5-0.5B-Inst. Van. 5.72 15.41 21.39 26.04 31.44
Qwen2.5-0.5B-Inst. AWQ 7.53 17.29 22.43 24.99 27.75
Qwen2.5-0.5B-Inst. GPTQ 7.17 17.35 22.44 24.99 28.05
Qwen2.5-1.5B-Inst. Van. 4.38 10.42 20.30 25.59 35.71
Qwen2.5-1.5B-Inst. AWQ 4.89 16.57 21.37 25.31 31.48
Qwen2.5-1.5B-Inst. GPTQ 4.41 14.22 21.03 23.82 36.53
Qwen2.5-3B-Inst. Van. 3.39 12.20 19.03 25.73 39.65
Qwen2.5-3B-Inst. AWQ 3.30 12.21 20.12 25.86 38.50
Qwen2.5-3B-Inst. GPTQ 3.16 15.27 19.52 26.50 35.55
Qwen2.5-7B-Inst. Van. 3.42 12.45 19.07 26.01 39.05
Qwen2.5-7B-Inst. AWQ 3.34 12.02 18.75 25.79 40.10
Qwen2.5-7B-Inst. GPTQ 3.38 12.00 19.31 25.19 40.12
Notes: Van. denotes the vanilla full-precision baseline; Inst. is an abbreviation of Instruct.
From the subject-wise analysis in Table 7, we observe a clear and consistent trend across all model scales: quantization-induced degradation is not uniformly distributed over mathematical domains. Subfields that involve multi-step symbolic manipulation, such as Intermediate Algebra, Precalculus and more advanced algebraic transformations, show noticeably larger performance drops for both AWQ and GPTQ. In contrast, domains that rely more on direct recall or simpler numerical reasoning remain comparatively stable. This pattern suggests that low-bit perturbations disproportionately affect tasks that require long-range dependency tracking and precise arithmetic transformations, which is consistent with the step-level error analysis in the main paper.
From the difficulty-level analysis in Table 8, we see a monotonic increase in degradation as problem difficulty grows. Level 1 and Level 2 questions exhibit only marginal changes after quantization, whereas degradation becomes much more pronounced for Levels 3 to 5. For Level 5 in particular, the gap between full-precision and quantized models can exceed 6–10 percentage points even for larger models.
These findings provide empirical evidence for the following points:
  • Conceptually demanding and algebraically heavy subfields of MATH are especially vulnerable to precision reduction.
  • Higher-difficulty problems, which require deeper chains of reasoning, tend to accumulate quantization noise and errors more severely.
Overall, this analysis further supports our claim that post-training quantization disproportionately affects mathematical reasoning compared with general language understanding. We include these tables and observations in the appendix for completeness.